OpenStudy (anonymous):

Which set of integers is a Pythagorean triple? A.10, 24, 25 B.9, 12, 21 C.8, 15, 23 D.6, 8, 10 not good with these at all. pplleeeaaassseee

OpenStudy (ashleyisakitty):

a^2+b^2=c^2. which two numbers add up to be the third?

OpenStudy (anonymous):

i dont understand.

OpenStudy (ashleyisakitty):

It's asking which number set is comparable to pythag. Pythag is a^2+b^2=c^2. So for the first one, 10 would be a, 24 would be c, and 25 b. Lets plug it in and see if it fits. 10+24=24. c=25 and not 24, so we know A is not the correct answer.

OpenStudy (ashleyisakitty):

Opps! I mean to say, 24 would be b, and 25 c. My bad!

OpenStudy (anonymous):

oh ok. i kinda understand. its kinda hard for me to understand some things. so it is kinda confusing me.

OpenStudy (ashleyisakitty):

Its okay :) Which one do you think it is?

OpenStudy (anonymous):

i kinda think it is b?

OpenStudy (anonymous):

please tell me if im wrong.

OpenStudy (ashleyisakitty):

You're forgetting to square the numbers :)

OpenStudy (ashleyisakitty):

9^2+12^2=225. 21^2=441

OpenStudy (anonymous):

so its not b?

OpenStudy (anonymous):

ookkk. i got that but i dont understand how to figure it out. can you give me the problem or something to help me better.

OpenStudy (ashleyisakitty):

no. 6^2=36 8^2=64. 36+64=100 10^2=100. therefore D) is the correct answer

OpenStudy (anonymous):

is it d?

OpenStudy (anonymous):

yay. thanks.

OpenStudy (ashleyisakitty):

you're welcome :)